How do u write the standard form equation of a graph with a slope of 7 and that goes through the point (1,2)
Answer by jim_thompson5910(35256)

y = mx+b

y = 7x+b

2 = 7(1)+b

2 = 7+b

2-7 = b

-5 = b

b = -5

y = mx+b

y = 7x - 5

y - 7x = -5

-7x + y = -5

7x - y = 5

So the answer is 7x - y = 5